### v1

@@ Abstract
Deep neural networks (DNNs) have exhibited substantial success in the field of Natural Language Processing (NLP). As these systems are increasingly integrated into real-world applications, ensuring their safety and reliability becomes a primary concern. There are safety critical contexts where such models must be robust to variability or attack, and give guarantees over their output. Computer Vision had pioneered the use of formal verification for neural networks for such scenarios and developed common verification standards and pipelines. In contrast, NLP verification methods have only recently appeared in the literature. While presenting sophisticated algorithms on their own right, these papers have not yet crystallised into a common methodology, they are often light on the pragmatical issues of NLP verification, and the area remains fragmented.
In this paper, we make an attempt to distil and evaluate general components of an NLP verification pipeline, that emerges from the progress in the field to date. Our contributions are two-fold.
Firstly, we give a general (i.e. algorithm-independent) characterisation of verifiable subspaces that result from embedding sentences into continuous spaces. We identify, and give an effective method to deal with, the technical challenge of semantic generalisability of verified subspaces; and propose it as a standard metric in the NLP verification pipelines (alongside with the standard metrics of model accuracy and model verifiability).
Secondly, we propose a general methodology to analyse the effect of the embedding gap – a problem that refers to the discrepancy between verification of geometric subpspaces on the one hand, and semantic meaning of sentences which the geometric subspaces are supposed to represent, on the other hand. In extreme cases, poor choices in embedding of sentences may invalidate verification results. We propose a number of practical NLP methods that can help to identify the effects of the embedding gap; and in particular we propose the metric of falsifiability of semantic subpspaces that we propose as another fundamental metric to be reported as part of the NLP verification pipeline.
We believe that together these general principles pave the way towards a more consolidated and effective development of this new domain.
0.1 Introduction
@@
Similarly, as such systems become widespread in their use, it may be desirable to have guarantees on queries concerning safety critical domains, for example when the user asks for medical advice. Research has shown that users tend to attribute undue expertise to systems [2, 34] potentially causing real world harm [12] (e.g. ‘Is it safe to take these painkillers with a glass of wine?’). However, a question remains on how to ensure that NLP systems can give formally guaranteed outputs, particularly for scenarios that require maximum control over the output.
Formal verification of neural networks.
@@
However, existing verification approaches primarily focus on computer vision (CV) tasks, where images are seen as vectors in a continuous space and every point in the space corresponds to a valid image. In contrast, sentences in NLP form a discrete domain [^2] In this paper, we work with textual representations of sentences. Raw audio input can be seen as continuous, but this is out of scope of this paper., making it challenging to apply traditional verification techniques effectively. In particular, taking an NLP dataset $\mathcal{Y}$ to be a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $\mathbb{E}$ is a function that maps a sentence to a vector in ${}^{m}$. The resulting vector space is called the embedding space. Due to discrete nature of the set $\mathcal{Y}$, the reverse of the embedding function $\mathbb{E}^{-1}:^{m}\rightarrow\mathcal{Y}$ is not total or undefined. This problem is known as the “problem of the embedding gap". Sometimes, one uses the term to more generally refer to any discrepancies that $\mathbb{E}$ introduces, for example, when it maps dissimilar sentences close in ${}^{m}$. We use the term in both mathematical and NLP sense.
@@
Mathematically, the general (geometric) “DNN robustness verification" approach of defining and verifying subspaces of ${}^{m}$ should work, and some prior works exploit this fact. However, pragmatically, because of the embedding gap only a tiny fraction of vectors contained in the verified subspaces maps back to valid sentences. When a verified subspace contains no or very few sentence embeddings, we say that verified subspace has low generalisability. Low generalisability may render verification efforts obsolete for practical applications.
@@
From the NLP perspective, there are other, more subtle, examples when the embedding gap can manifest. Suppose we succeeded in verifying a DNN robust on some subspace of ${}^{m}$. Consider an example when the subspace contains sentences that are semantically similar to the sentence: i really like too chat to a human. are you one?. Verification gives us guarantees that the DNN will always identify these sentences as questions about human/robot identity. But suppose the embedding function $\mathbb{E}$ wrongly embedded sentences belonging to an opposite class into this subspace. For example, an LLM Vicuna generates the following sentence as a rephrase of the previous one: Do you take pleasure in having a conversation with someone?. Suppose our verified subspace contained an embedding of this sentence, too, and thus our verified DNN identifies this second sentence to belong to the the same class as the first one. However, the second sentence is not a question about human/robot identity of the agent! When we can find such an example, we say that it falsifies the verification guarantee for the subspace it is contained in. Alternatively, we say that the subpspace is falsifiable.
@@
Our main aim is to provide a general and principled verification methodology that bridges the embedding gap when possible; and gives precise metrics to evaluate and report its effects in any case. The contributions split into two main groups, depending whether the embedding gap is approached from mathematical or NLP perspective.
@@
Contributions Part 1: Characterisation of verifiable subspaces and NLP verification pipeline.
@@
An alternative method to purely geometric approach that suffers from the embedding gap is to construct subspaces of the embedding space based on the semantic perturbations of sentences [54, 50, 146]. Concretely, the idea is to form each $\mathcal{S}_{i}$ by embedding a sentence $i$ and its $n$ semantic perturbations into the real vector space and enclosing them inside some geometric shape. Ideally, such a shape should be given by a convex hull around these $n+1$ embedded sentences, however calculating convex hulls with sufficient precision is computationally infeasible for high number of dimensions. Thus, simpler shapes, such as hyper-cubes and hyper-rectangles are used in the literature. We propose a novel refinement of these ideas, by including the method of a hyper-rectangle rotation in order to increase the shape precision (see Figure 1). We will call the resulting shapes semantic subspaces (in contrast to those obtained purely geometrically).
A few questions have been left un-answered in the previous work. Firstly, because generalisability of the verified subspaces is not reported in the literature, we cannot know whether the prior semantically-informed approaches are better in that respect than purely geometric methods. If they are better in both verifiability and generalisability, it is unclear whether the improvement should be attributed to:
@@
This paper provides a strong argument for including generalisability as a standard metric in reporting NLP verification results in the future. Through a series of experiments, we confirm that verified semantic subspaces are more verifiable and more generalisable than their geometric counterparts. Moreover, by comparing the volumes of the obtained verified semantic and geometric subspaces, we show that the improvement is partly due to finding an optimal size of subspaces (for the given embedding space), and partly due to improvement in shape precision.
The second group of unresolved questions concerns robust training regimes in NLP verification that is used as means of improving verifiability of subspaces in prior works [54, 50, 146]. It was not clear what made robust training successful:
  • was it because additional examples generally improved the precision of the decision boundary? (in which case data set augmentation would have a similar effect);
@@
We argue that, together with the new generalisability metric, this approach opens the way for more principled evaluation of performance of NLP verification methods that accounts for the effects of the embedding gap; and generation of more transparent NLP verification benchmarks. We implement a tool that generates NLP verification benchmarks based on the above choices. This paper is the first to use a complete SMT-based verifier (namely Marabou [131]) for NLP verification.
@@
We test the theoretical results by suggesting an NLP verification pipeline, a general methodology that starts with NLP analysis of the dataset and obtaining semantically similar perturbations that together characterise the semantic meaning of a sentence; proceeds with embedding of the sentences into the real vector space and defining semantic subspaces around embeddings of semantically similar sentences; and culminates with using these subspaces for both training and verification. This clear division into stages allows us to formulate practical NLP methods for minimising the effects of the embedding gap. In particular, we show that the quality of the generated sentence perturbations maybe improved through the use of human evaluation, cosine similarity and ROUGE-N. We introduce the falsifiability metric as an effective practical way to measure the quality of the embedding functions. Through a detailed case study, we show how geometric and NLP intuitions can be put together at work towards obtaining DNNs that are more verifiable over better generalisable and less falsifiable semantic subspaces. Perhaps more importantly, the proposed methodology opens the way for transparency in reporting NLP verification results, – something that this domain will benefit from if it reaches the stage of practical deplyoment of NLP verification pipelines.
Paper Outline.
@@
From here, the paper proceeds as follows. Section 0.2 gives an extensive literature review encompassing DNN verification methods generally, and NLP verification methods in particular. The section culminates with distilling a common “NLP verification pipeline" encompassing the existing literature. Based on the understanding of major components of the pipeline, the rest of the paper focuses on improving understanding or implementation of its components. Section 0.3 formally defines the components of the pipeline in a general mathematical notation, which abstracts away from particular choices of sentence perturbation, sentence embedding, training and verification algorithms. The central notion the section introduces is that of geometric and semantic subpspaces. The next Section 0.4 makes full use of this general definition, and shows that semantic subpspaces play a pivotal role in improving verification and training of DNNs in NLP. This section formally defines the generalisability metric and considers the problem of generalisability-verifiability trade-off. Through thorough empirical evaluation, it shows that a principled approach to defining semantic subpaces can help to improve both generalisability and verifiability of DNNs, thus reducing the effects of the trade-off. The final Section 0.5 further tests the NLP verification pipelines using state-of-the-art NLP tools, and analyses the effects of the embedding gap from the NLP persective, in particular it introduces a method of measuring falsifiability of semantic subpaces and reporting this metric alongside verifiability and generalisability. Section 0.6 concludes the paper and discusses future work.
0.2 Related Work
0.2.1 DNN Verification
@@
Formal verification is an active field across several domains including hardware [64, 97], software languages [58], network protocols [87] and many more [129], however it was only recently that this became applicable to the field of machine learning [59]. Several verifiers have been popular in DNNs verification and competitions [5, 6, 81, 108]. We can divide them into 2 main categories: complete and incomplete verifiers. When the verification approach guarantees that, if a query fails, then the query is false and it produces a counter-example to prove it, then we call it complete verification, otherwise we call it incomplete verification. Furthermore, while complete verifiers are always deterministic, incomplete verifiers may also be probabilistic. Unlike deterministic verification, probabilistic verification is guaranteed to output ‘not verified’ with a certain probability (e.g., 99.9%).
Complete verifiers
can be based on Satisfiability Modulo Theories (SMT), Mixed-Integer Linear Programming (MILP) or Branch-and-Bound (Bab). SMT-based verification [99, 59, 131] is built upon the observation that feed-forward neural networks are defined by the sequential composition of affine trasformations and ReLU operations. Both these transformations and operations can be encoded by a conjunction of linear inequalities, thus general-purpose SMT solvers can be directly applied to solve the satisfiability problem, yielding a solution to complete verification. A state-of-the-art SMT-based tool is Marabou [131], which answers queries about neural networks and their properties in the form of constraint satisfaction problems. Marabou takes as input networks with piece-wise linear activation functions and with fully connected topology. It first applies multiple pre-processing steps to infer bounds for each node in the network. Next it applies a combination of Simplex [31] search over linear constraints with SMT techniques directing the search over non-linear constraints.
MILP-based approaches [23, 82, 114] encode the verification problem as a mixed-integer linear programming problem, in which the constraints are linear inequalities and the objective is represented by a linear function. Differently from linear programming (LP), in MILP it is possible to constrain some variables to take only integer values instead of real numbers, allowing the constraints to encode the non-linear ReLU operations. Thus, the verification problem can be precisely encoded as an MILP problem. A representative tool for this category is ERAN [109], which is mainly based on abstract interpretation (see ‘Incomplete verifiers’ below ) but can also leverage the efficient MILP solver GUROBI [48]. ERAN combines abstract domains with custom multi-neuron relaxations to support fully-connected, convolutional, and residual networks with ReLU, Sigmoid, Tanh, and Maxpool activations. Both these methods suffer from scalability, which is their main limitation, but their strength is that they precisely encode the constraint.
BaB-based verification [43, 19, 18, 41, 57, 118, 140] relies on the piecewise-linear property of DNNs: since each ReLU neuron outputs ReLU($x$) = max{$x$, 00}, it is always locally linear within some region around input $x$. Furthermore, since feed-forward ReLU networks are the composition of these piecewise linear neurons and (linear) affine transformations, the output is locally linear w.r.t. input $x$. This property is formally stated and proved in [57] and it serves as the foundation for BaB verification. A BaB verification approach, as the name suggests, consists of two parts: branching and bounding. It first applies incomplete verification to derive a lower bound and an upper bound, then, if the lower bound is positive it terminates with ‘verified’, else, if the upper bound is non-positive it terminates with ‘not verified’ (bounding). Otherwise, the approach recursively chooses a neuron to split into two branches (branching), resulting in two linear constraints. Then bounding is applied to both constraints and if both are satisfied the verification terminates, otherwise the other neurons are split recursively. When all neurons are split, the branch will contain only linear constraints, and thus the approach applies linear programming to compute the precise constraint and verify the branch. Multi-Neuron Guided Branch-and-Bound (MN-BaB) [41] is a state-of-the-art neural network verifier that builds on the tight multi-neuron constraints proposed in PRIMA [95] and leverages these constraints within a BaB framework to yield an efficient, GPU based dual solver. Another state-of-the-art tool is $\alpha,\beta$-CROWN [133, 118], a neural network verifier based on an efficient linear bound propagation framework and branch-and-bound. It can be accelerated efficiently on GPUs and can scale to relatively large convolutional networks (e.g., $10^{7}$ parameters). It also supports a wide range of neural network architectures (e.g., CNN, ResNet, and various activation functions). BaB-based methods are more scalable than solver-based approaches, however they introduce a level of abstraction and sacrifice precision in favor of scalability. For example GCP-CROWN [140] extracts convex constraints from MILP solvers and integrates them in linear inequality propagation, which can be viewed as leveraging multi-neuron relaxations in branch-and-bound complete verification.
Deterministic incomplete verifiers
@@
mainly use linear relaxations on ReLU neurons, resulting in an over-approximation of the initial constraint. In general, they define a lower bound and an upper bound of the output of each ReLU neuron as linear constraints, which define a region called ReLU polytope that gets propagated through the network. A predominant approach is the use of interval bound propagation (IBP) [128, 46, 83, 88]. The strength of IBP-based methods lies in their efficiency; they are faster than alternative approaches and demonstrate superior scalability. However, their primary limitation lies in the inherently loose bounds they produce [46]. This drawback becomes particularly pronounced in the case of deeper neural networks, typically those with more than 10 layers [73], where they cannot certify non-trivial robustness due to the amplification of over-approximation. Other methods that are less efficient but produce tighter bounds are based on polyhedra abstraction, such as CROWN [141] and DeepPoly [111], or based on multi-neuron relaxation, such as PRIMA [93]. One of the most mature tool in this category is ERAN [109], which can be used for complete verification, but its main purpose is deterministic incomplete verification through abstract interpretation (DeepPoly) and multi-neuron relaxation (PRIMA).
Probabilistic incomplete verification
@@
approaches add random noise to smooth models, and then derive certified robustness for these smoothed models. This field is commonly referred to as Randomised Smoothing, given that these approaches provide probabilistic guarantees of robustness, and all current probabilistic verification techniques are tailored for smoothed models [67, 71, 38, 138, 103, 89]. Given that this work focuses on deterministic approaches, here we only report the existence of this line of work without going into details.
@@
0.2.2 Robust Training
Verifying DNNs poses significant challenges if they are not appositely trained. The fundamental issue lies in the failure of DNNs, including even sophisticated models, to meet essential verification properties, such as robustness [22]. To enhance robustness, various training methodologies have been proposed. It is noteworthy that, although robust training by projected gradient descent [45, 85, 62] predates verification, contemporary approaches are often related to, or derived from, the corresponding verification methods by optimizing verification-inspired regularization terms or injecting specific data augmentation during training. In practice, after robust training, the model usually achieves higher certified robustness and is more likely to satisfy the desired verification properties [22]. Thus, robust training is a strong complement to robustness verification approaches.
Robust training
techniques can be classified into several large groups:

### v3

@@ Abstract
Machine Learning (ML) has exhibited substantial success in the field of Natural Language Processing (NLP). For example large language models (LLMs) have empirically proven to be capable of producing text of high complexity and cohesion. However, at the same time, they are prone to inaccuracies and hallucinations. As these systems are increasingly integrated into real-world applications, ensuring their safety and reliability becomes a primary concern. There are safety critical contexts where such models must be robust to variability or attack, and give guarantees over their output. Computer Vision had pioneered the use of formal verification of neural networks for such scenarios and developed common verification standards and pipelines, leveraging precise formal reasoning about geometric properties of data manifolds. In contrast, NLP verification methods have only recently appeared in the literature. While presenting sophisticated algorithms in their own right, these papers have not yet crystallised into a common methodology. They are often light on the pragmatical issues of NLP verification, and the area remains fragmented.
In this paper, we attempt to distil and evaluate general components of an NLP verification pipeline, that emerges from the progress in the field to date. Our contributions are two-fold. Firstly, we propose a general methodology to analyse the effect of the embedding gap – a problem that refers to the discrepancy between verification of geometric subspaces, and the semantic meaning of sentences which the geometric subspaces are supposed to represent. We propose a number of practical NLP methods that can help to quantify the effects of the embedding gap. Secondly, we give a general method for training and verification of neural networks that leverages a more precise geometric estimation of semantic similarity of sentences in the embedding space and helps to overcome the effects of the embedding gap in practice.
keywords:
Neural Networks, Verification, Natural Language Processing, Robustness, Adversarial Training, Machine Learning
{Frontmatter}
[…]
Casadio M. et al.
[…]
Casadio M. et al.
1 Introduction
@@
Similarly, as such systems become widespread in their use, it may be desirable to have guarantees on queries concerning safety critical domains, for example when the user asks for medical advice. Research has shown that users tend to attribute undue expertise to NLP systems [12, 7] potentially causing real world harm [13] (e.g. ‘Is it safe to take these painkillers with a glass of wine?’). However, a question remains on how to ensure that NLP systems can give formally guaranteed outputs, particularly for scenarios that require maximum control over the output.
@@
Existing verification approaches primarily focus on computer vision (CV) tasks, where images are seen as vectors in a continuous space and every point in the space corresponds to a valid image. In contrast, sentences in NLP form a discrete domain [^2] In this paper, we work with textual representations of sentences. Raw audio input can be seen as continuous, but this is out of scope of this paper., making it challenging to apply traditional verification techniques effectively. In particular, taking an NLP dataset $\mathcal{Y}$ to be a set of sentences $s_{1},\ldots,s_{q}$ written in natural language, an embedding $E$ is a function that maps a sentence to a vector in m. The resulting vector space is called the embedding space. Due to discrete nature of the set $\mathcal{Y}$, the reverse of the embedding function $E^{-1}:^{m}\rightarrow\mathcal{Y}$ is undefined for some elements of m. This problem is known as the “problem of the embedding gap”. Sometimes, one uses the term to more generally refer to any discrepancies that $E$ introduces, for example, when it maps dissimilar sentences close in m. We use the term in both mathematical and NLP sense.
@@
Mathematically, the general (geometric) “DNN robustness verification” approach of defining and verifying subspaces of m should work, and some prior works exploit this fact. However, pragmatically, because of the embedding gap, usually only a tiny fraction of vectors contained in the verified subspaces map back to valid sentences. When a verified subspace contains no or very few sentence embeddings, we say that verified subspace has low generalisability. Low generalisability may render verification efforts ineffective for practical applications.
@@
From the NLP perspective, there are other, more subtle, examples where the embedding gap can manifest. Consider an example of a subspace containing sentences that are semantically similar to the sentence: ‘i really like too chat to a human. are you one?’. Suppose we succeed in verifying a DNN to be robust on this subspace. This provides a guarantee that the DNN will always identify sentences in this subspace as questions about human/robot identity. But suppose the embedding function $E$ wrongly embeds sentences belonging to an opposite class into this subspace. For example, the LLM Vicuna [14] generates the following sentence as a rephrasing of the previous one: Do you take pleasure in having a conversation with someone?. Suppose our verified subspace contained an embedding of this sentence too, and thus our verified DNN identifies this second sentence to belong to the same class as the first one. However, the second sentence is not a question about human/robot identity of the agent! When we can find such an example, we say that the verified subspace is prone to embedding errors.
Robustness verification in NLP is particularly susceptible to this problem, because we cannot cross the embedding gap in the opposite direction as the embedding function is not invertible. This means it is difficult for humans to understand what sort of sentences are captured by a given subspace.
@@
Our main aim is to provide a general and principled verification methodology that bridges the embedding gap when possible; and gives precise metrics to evaluate and report its effects in any case. The contributions split into two main groups, depending on whether the embedding gap is approached from mathematical or NLP perspective.
Contributions Part 1: Characterisation of Verifiable Subspaces and general
@@
An alternative method to the purely geometric approach is to construct subspaces of the embedding space based on the semantic perturbations of sentences (first attempts to do this appeared in [16, 17, 18, 19]). Concretely, the idea is to form each $\mathcal{S}_{i}$ by embedding a sentence $s$ and $n$ semantic perturbations of $s$ into the real vector space and enclosing them inside some geometric shape. Ideally, this shape should be the convex hull around the $n+1$ embedded sentences (see Figure 1), however calculating convex hulls with sufficient precision is computationally infeasible for high number of dimensions. Thus, simpler shapes, such as hyper-cubes and hyper-rectangles are used in the literature. We propose a novel refinement of these ideas, by including the method of a hyper-rectangle rotation in order to increase the shape precision (see Figure 1). We will call the resulting shapes semantic subspaces (in contrast to those obtained purely geometrically).
A few questions have been left unanswered in the previous work [16, 17, 18, 19]. Firstly, because generalisability of the verified subspaces is not reported in the literature, we cannot know whether the prior semantically-informed approaches are better in that respect than purely geometric methods. If they are better in both verifiability and generalisability, it is unclear whether the improvement should be attributed to:
@@
Through a series of experiments, we confirm that semantic subspaces are more verifiable and more generalisable than their geometric counterparts. Moreover, by comparing the volumes of the obtained verified semantic and geometric subspaces, we show that the improvement is partly due to finding an optimal size of subspaces (for the given embedding space), and partly due to improvement in shape precision.
The second group of unresolved questions concerns robust training regimes in NLP verification that is used as means of improving verifiability of subspaces in prior works [16, 17, 18, 19]. It was not clear what made robust training successful:
  • was it because additional examples generally improved the precision of the decision boundary (in which case dataset augmentation would have a similar effect);
@@
We argue that this approach opens the way for more principled NLP verification methods that reduces the effects of the embedding gap; and generation of more transparent NLP verification benchmarks. We implement a tool ANTONIO that generates NLP verification benchmarks based on the above choices. This paper is the first to use a complete SMT-based verifier (namely Marabou [22]) for NLP verification.
@@
We test the theoretical results by suggesting an NLP verification pipeline, a general methodology that starts with NLP analysis of the dataset and obtaining semantically similar perturbations that together characterise the semantic meaning of a sentence; proceeds with embedding of the sentences into the real vector space and defining semantic subspaces around embeddings of semantically similar sentences; and culminates with using these subspaces for both training and verification. This clear division into stages allows us to formulate practical NLP methods for minimising the effects of the embedding gap. In particular, we show that the quality of the generated sentence perturbations maybe improved through the use of human evaluation, cosine similarity and ROUGE-N. We introduce the novel embedding error metric as an effective practical way to measure the quality of the embedding functions. Through a detailed case study, we show how geometric and NLP intuitions can be put at work towards obtaining DNNs that are more verifiable over better generalisable and less prone to embedding errors semantic subspaces. Perhaps more importantly, the proposed methodology opens the way for transparency in reporting NLP verification results, – something that this domain will benefit from if it reaches the stage of practical deployment of NLP verification pipelines.
@@
Paper Outline. From here, the paper proceeds as follows. Section 2 gives an extensive literature review encompassing DNN verification methods generally, and NLP verification methods in particular. The section culminates with distilling a common “NLP verification pipeline” encompassing the existing literature. Based on the understanding of major components of the pipeline, the rest of the paper focuses on improving understanding or implementation of its components. Section 3 formally defines the components of the pipeline in a general mathematical notation, which abstracts away from particular choices of sentence perturbation, sentence embedding, training and verification algorithms. The central notion the section introduces is that of geometric and semantic subspaces. The next Section 4 makes full use of this general definition, and shows that semantic subspaces play a pivotal role in improving verification and training of DNNs in NLP. This section formally defines the generalisability metric and considers the problem of generalisability-verifiability trade-off. Through thorough empirical evaluation, it shows that a principled approach to defining semantic subspaces can help to improve both generalisability and verifiability of DNNs, thus reducing the effects of the trade-off. The final Section 5 further tests the NLP verification pipelines using state-of-the-art NLP tools, and analyses the effects of the embedding gap from the NLP perspective, in particular it introduces a method of measuring the embedding error and reporting this metric alongside verifiability and generalisability. Section 6 concludes the paper and discusses future work.
2 Related Work
2.1 DNN Verification
@@
Formal verification is an active field across several domains including hardware [23, 24], software [25], network protocols [26] and many more [27]. However, it was only recently that this became applicable to the field of machine learning [28]. An input query to a verifier consists of a subspace within the embedding space and a target subspace of outputs, typically a target output class. The verifier then returns either true, false or unknown. True indicates that there exists an input within the given input subspace whose output falls within the given output subspace, often accompanied by an example of such input. False indicates that no such input exists. Several verifiers are popular in DNN verification and competitions [29, 30, 31, 32]. We can divide them into 2 main categories: complete verifiers which return true/false and incomplete verifiers which return true/unknown. While complete verifiers are always deterministic, incomplete verifiers may be probabilistic. Unlike deterministic verification, probabilistic verification is not sound and a verifier may incorrectly output true with a very low probability (typically 0.01%).
Complete Verification based on Linear Programming & Satisfiability Modulo Theories (SMT) solving. Generally, SMT solving is a group of methods for determining the satisfiability of logical formulas with respect to underlying mathematical theories such as real arithmetic, bit-vectors, or arrays [33]. These methods extend traditional satisfiability (SAT) solving by incorporating domain-specific reasoning, making them particularly useful for verifying complex systems. In the context of neural network verification, SMT solvers encode network behaviours and safety properties as logical constraints, enabling rigorous checks for violations of specifications [34]. When the activation functions are piecewise linear (e.g. ReLU), the DNN can be encoded by conjunctions and disjunctions of linear inequalities and thus linear programming algorithms can be directly applied to solve the satisfiability problem. A state-of-the-art tool is Marabou [22], which answers queries about neural networks and their properties in the form of constraint satisfaction problems. Marabou takes the network as input and first applies multiple pre-processing steps to infer bounds for each node in the network. It applies the algorithm ReLUplex [28], a combination of Simplex [35] search over linear constraints, modified to work for networks with piece-wise linear activation functions. With time, Marabou grew into a complex prover with multiple heuristics supplementing the original ReLUplex algorithm [22], for example it now includes mixed-integer linear programming (MILP) [36] and abstract interpretation based algorithms which we survey below. MILP-based approaches [37, 38, 39] encode the verification problem as a mixed-integer linear programming problem, in which the constraints are linear inequalities and the objective is represented by a linear function. Thus, the DNN verification problem can be precisely encoded as a MILP problem. For example, ERAN [40] combines abstract interpretation with the MILP solver GUROBI [41]. By the time Branch and Bound (BaB) methodologies are introduced later, it becomes evident that the verification community has effectively consolidated diverse approaches into a unified taxonomy. Modern verifiers, such as $\alpha\beta$-CROWN [42, 43], take full advantage of this combination and effectively balance efficiency with precision.
@@
Incomplete Verification based on Abstract Interpretation takes inspiration from the domain of abstract interpretation, and mainly uses linear relaxations on ReLU neurons, resulting in an over-approximation of the initial constraint. Abstract interpretation was first developed by Cousot and Cousot [44] in 1977. It formalises the idea of abstraction of mathematical structures, in particular those involved in the specification of properties and proof methods of computer systems [45] and it has since been used in many applications [46]. Specifically, for DNN verification, this technique can model the behaviour of a network using an abstract domain that captures the possible range of values the network can output for a given input. Abstract interpretation-based verifiers can define a lower bound and an upper bound of the output of each ReLU neuron as linear constraints, which define a region called ReLU polytope that gets propagated through the network. To propagate the bounds, one can use interval bound propagation (IBP) [47, 48, 49, 50]. The strength of IBP-based methods lies in their efficiency; they are faster than alternative approaches and demonstrate superior scalability. However, their primary limitation lies in the inherently loose bounds they produce [48]. This drawback becomes particularly pronounced in the case of deeper neural networks, typically those with more than 10 layers [51], where they cannot certify non-trivial robustness properties. Other methods that are less efficient but produce tighter bounds are based on polyhedra abstraction, such as CROWN [52] and DeepPoly [53], or based on multi-neuron relaxation, such as PRIMA [54]. An abstract interpretation tool CORA [55], uses polyhedral abstractions and reachability analysis for formal verification of neural networks. It integrates various set representations, such as zonotopes, and algorithms to compute reachable sets for both continuous and hybrid systems, providing tighter bounds in verification tasks. Another mature tool in this category is ERAN [40], which uses abstract domains (DeepPoly) with custom multi-neuron relaxations (PRIMA) to support fully-connected, convolutional, and residual networks with ReLU, Sigmoid, Tanh, and Maxpool activations. Note that, having lost completeness, they can work with a more general class of neural networks (e.g. neural networks with non linear layers).
Modern Neural Network Verifiers. Modern verifiers are complex tools that take advantage of a combination of complete and incomplete methods as well as additional heuristics. The term Branch and Bound (BaB) [56, 57, 58, 59, 60, 43, 61] often refers to the method that relies on the piecewise linear property of DNNs: since each ReLU neuron outputs ReLU($x$) = max{$x$,00} is piecewise linear, we can consider its two linear pieces $x\geq 0$, $x\leq 0$ separately. A BaB verification approach, as the name suggests, consists of two parts: branching and bounding. It first derives a lower bound and an upper bound, then, if the lower bound is positive it terminates with ‘verified’, else, if the upper bound is non-positive it terminates with ‘not verified’ (bounding). Otherwise, the approach recursively chooses a neuron to split into two branches (branching), resulting in two linear constraints. Then bounding is applied to both constraints and if both are satisfied the verification terminates, otherwise the other neurons are split recursively. When all neurons are split, the branch will contain only linear constraints, and thus the approach applies linear programming to compute the constraint and verify the branch. It is important to note that BaB approaches themselves are neither inherently complete nor incomplete. BaB is an algorithm for splitting problems into sub-problems and requires a solver to resolve the linear constraints. The completeness of the verification depends on the combination of BaB and the solver used. Multi-Neuron Guided Branch-and-Bound (MN-BaB) [59] is a state-of-the-art neural network verifier that builds on the tight multi-neuron constraints proposed in PRIMA [62] and leverages these constraints within a BaB framework to yield an efficient, GPU based dual solver. Another state-of-the-art tool is $\alpha\beta$-CROWN [42, 43], a neural network verifier based on an efficient linear bound propagation framework and branch-and-bound. It can be accelerated efficiently on GPUs and can scale to relatively large convolutional networks (e.g., $10^{7}$ parameters). It also supports a wide range of neural network architectures (e.g., CNN, ResNet, and various activation functions).
@@
Probabilistic Incomplete Verification approaches add random noise to models to smooth them, and then derive certified robustness for these smoothed models. This field is commonly referred to as Randomised Smoothing, given that these approaches provide probabilistic guarantees of robustness, and all current probabilistic verification techniques are tailored for smoothed models [63, 64, 65, 66, 67, 68]. Given that our work focuses on deterministic approaches, here we only report the existence of this line of work without going into details.
@@
2.2 Geometric Representations in DNN Verification
[…]
2.3 Robust Training
Verifying DNNs poses significant challenges if they are not appropriately trained. The fundamental issue lies in the failure of DNNs, including even sophisticated models, to meet essential verification properties, such as robustness [70]. To enhance robustness, various training methodologies have been proposed. It is noteworthy that, although robust training by projected gradient descent [71, 20, 72] predates verification, contemporary approaches are often related to, or derived from, the corresponding verification methods by optimizing verification-inspired regularization terms or injecting specific data augmentation during training. In practice, after robust training, the model usually achieves higher certified robustness and is more likely to satisfy the desired verification properties [70]. Thus, robust training is a strong complement to robustness verification approaches.
Robust training techniques can be classified into several large groups: